# Offline-online approximation of multiscale eigenvalue problems with random defects

### v1

@@ 5.1.2 Two-dimensional experiments
Figure 4 illustrates two-dimensional results for the random checkerboard (left) and random erasure (right) shown in Figure 1 in Section 1. The parameters are set to $d=2$, $\varepsilon=2^{-7}$ and $h=2^{-8}$.
Again, our OLOD method shows a very promising performance with a RMSE of a few percent for reasonable defect probabilities. For instance, the random checkerboard in Figure 4(a) shows RMSEs less than $10\%$ for $H=2^{-6}$ and a defect probability up to $10\%$. In the random erasure, Figure 4(b), RMSEs are even lower, more precisely for $H=2^{-6}$, we observe less than $2\%$ error for material defects up to $10\%$.
@@
Figure 4: The two-dimensional random checkerboard for $\varepsilon=2^{-7}$ with a sample size $200$.
@@ 5.2 Experiments with the alternate offline-online strategy
In this section, we compare the accuracy of the OLOD and its alternate offline-online strategy (from Section 2.5) for the random checkerboard. The one- and two-dimensional comparisons are depicted in the Figures 5 and 6, respectively. The domain, spectral bounds and the LOD parameters are the same as in Section 5.1. We only compare the results for the smaller mesh sizes $H=2^{-5}$ and $H=2^{-6}$ to focus on the consistency error and reduce the influence of any discretization errors. In both dimensions the alternate method produces improved results for larger $p$ as expected. Particularly in two dimensions, we can observe an RMSE less than $\sim 6\%$ for about $10\%$ of material defects. For the one-dimensional case, as expected both methods produce better results with an RMSE of $\sim 4\%$ at a $10\%$ material error. However, for much smaller probabilities of material error in the $1$D case, slightly better results can be noticed for the OLOD with sum constraint one. We emphasize that this is no contradiction of our theoretical derivations as we only considered limits $\varepsilon\to 0$ there. As expected from the choice of $s$, the difference between the sum constraint one model and the alternate variant becomes more prominent with growing $p$. We emphasize that, although our derivations were based on one-dimensional heuristics, the two-dimensional results are very promising and underline the potential of our approach.
Figure 5: Comparison of offline-online strategy with sum constraint one and alternate version in one-dimension for $\varepsilon=2^{-7}$ over $200$ samples.
Figure 6: Comparison of offline-online strategy with sum constraint one and alternate version in two-dimension for $\varepsilon=2^{-7}$ over $200$ samples.
@@ Acknowledgments
This research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under project number 496556642. BV additionally acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.

### v2

@@ 5.1.2 Two-dimensional experiments
Figure 4 illustrates two-dimensional results for the random checkerboard (Figure 1(a)) and random erasure (Figure 1(b)) in Section 1.
Again, our OLOD method shows a very promising performance with a RMSE of a few percent for reasonable defect probabilities. For instance, the random checkerboard results in Figure 4(a) shows RMSEs less than $12\%$ for $H=2^{-5}$ and a defect probability up to $10\%$. In the random erasure, Figure 4(b), RMSEs are even lower, more precisely for $H=2^{-5}$, we observe less than $2\%$ error for material defects up to $10\%$.
@@
Figure 4: The OLOD (sum constraint one) results in two-dimensions.
@@ 5.2 Experiments with the alternate offline-online strategy
In this section, we compare the accuracy of the OLOD and its alternate offline-online strategy (from Section 2.5) for the random checkerboard. The one- and two-dimensional comparisons are depicted in the Figures 5 and 6, respectively. The $H_{uty}$ and $H_{s}$ denote the $H$-discretization corresponding to the sum constraint one and the alternate methods respectively. We only compare the results for the smaller mesh sizes $H=2^{-5}$ and $H=2^{-6}$ to focus on the consistency error and reduce the influence of any discretization errors. In both dimensions the alternate method produces improved results for larger $p$ as expected. Particularly in two dimensions, we can observe an RMSE less than $\sim 6\%$ for about $10\%$ of material defects. For the one-dimensional case, as expected both methods produce better results with an RMSE of $\sim 4\%$ at a $10\%$ material error. However, for much smaller probabilities of material error in the $1$D case, slightly better results can be noticed for the OLOD with sum constraint one. We emphasize that this is no contradiction of our theoretical derivations as we only considered limits $\varepsilon\to 0$ there. As expected from the choice of $s$, the difference between the sum constraint one model and the alternate variant becomes more prominent with growing $p$. We emphasize that, although our derivations were based on one-dimensional heuristics, the two-dimensional results are very promising and underline the potential of our approach. We note that in some practical scenarios in the context of metamaterials, defect probabilities are typically kept below $5\%$ (and sometimes even below $1\%$). Therefore, the defect probabilities considered in our experiments are already on the higher end and cover a broad range of realistic cases.
Figure 5: Comparison of the offline-online strategy with sum constraint one and alternate version in one-dimension.
Figure 6: Comparison of the offline-online strategy with sum constraint one and alternate version in two-dimension.
@@ Acknowledgments
This research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under project number 496556642. BV additionally acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. The authors would like to thank the anonymous reviewers for their constructive feedback and insightful comments, which helped improve the clarity and quality of this work.